# Beautiful Images, Toxic Words: Understanding and Addressing Offensive Text in Generated Images

### v3

@@ Abstract
State-of-the-art Diffusion Models (DMs) produce highly realistic images. While prior work has successfully mitigated Not Safe For Work (NSFW) content in the visual domain, we identify a novel threat: the generation of NSFW text embedded within images. This includes offensive language, such as insults, racial slurs, and sexually explicit terms, posing significant risks to users. We show that all state-of-the-art DMs (e.g., SD3, SDXL, Flux, DeepFloyd IF) are vulnerable to this issue. Through extensive experiments, we demonstrate that existing mitigation techniques, effective for visual content, fail to prevent harmful text generation while substantially degrading benign text generation. As an initial step toward addressing this threat, we introduce a novel fine-tuning strategy that targets only the text-generation layers in DMs. Therefore, we construct a safety fine-tuning dataset by pairing each NSFW prompt with two images: one with the NSFW term, and another where that term is replaced with a carefully crafted benign alternative while leaving the image unchanged otherwise. By training on this dataset, the model learns to avoid generating harmful text while preserving benign content and overall image quality. Finally, to advance research in the area, we release ToxicBench, an open-source benchmark for evaluating NSFW text generation in images. It includes our curated fine-tuning dataset, a set of harmful prompts, new evaluation metrics, and a pipeline that assesses both NSFW-ness and text and image quality. Our benchmark aims to guide future efforts in mitigating NSFW text generation in text-to-image models—contributing to their safe deployment. The benchmark is available at https://github.com/sprintml/ToxicBench.
Warning: This paper contains examples of offensive language, including insults, and sexual or explicit terms, used solely for research and analysis purposes.
@@ 1 Introduction
State-of-the-art Diffusion Models (DMs) (Esser et al., 2024; StabilityAI, 2023; Black Forest Labs, 2024), have revolutionized the creation of realistic, detailed, and aesthetically impressive content. Despite their capabilities, these models often raise ethical and safety concerns, as they can inadvertently generate Not Safe For Work (NSFW) content, such as depictions of violence or nudity (Qu et al., 2023; Rando et al., 2022; Yang et al., 2024a).
@@
To mitigate the generation of NSFW content, prior work has focused extensively on addressing such issues in the visual space. Beyond the development of powerful NSFW detectors (Berg, ; notAI tech, ), these efforts, which include modifying training data (Zong et al., 2024), adding safety-based loss functions (Poppi et al., 2025; Gandikota et al., 2023), and steering generation to safe subspaces (Schramowski et al., 2023), have shown promising results in reducing explicit or harmful visual scenes.
@@
However, as DMs have grown more powerful, they now go beyond visual generation. Instead, they also generate embedded text within those images, such as captions, signs, or artistic typography (Esser et al., 2024; Chen et al., 2023; StabilityAI, 2023; Black Forest Labs, 2024). This advancement introduces a new challenge: as we show in Figure 1, all prominent state-of-the-art DMs, such as SD3 (Esser et al., 2024), Flux (Black Forest Labs, 2024), DeepFloyd IF (StabilityAI, 2023) and SDXL (Podell et al., 2023), can inadvertently produce NSFW or offensive text, such as explicit language or slurs that can be deeply offensive to viewers and raise significant ethical concerns.
@@
We demonstrate that existing NSFW mitigation techniques (Gandikota et al., 2023; Poppi et al., 2025; Suau et al., 2024), while effective in addressing NSFW content in the visual or the language domain, are inadequate for handling embedded NSFW text in generated images without significantly degrading the models’ overall and (benign) text generation ability.
@@
Figure 1: Visual generative models output images with NSFW text. We evaluate the state-of-the-art DMs (SD3, DeepFloyd IF, FLUX, and SDXL) and observe that they easily generate NSFW text in the output images due to the lack of safety guardrails.
@@
As a first step toward addressing this threat, we introduce a novel method that performs lightweight fine-tuning on text-generation-relevant layers in DMs, previously identified by Staniszewski et al. (2025). By applying LoRA-based updates only to those layers, we enable efficient and focused mitigation. To supervise the intervention, we curate a safety fine-tuning dataset consisting of NSFW and benign prompt pairs that differ by a single word, where the harmful term is replaced with a carefully chosen benign counterpart. We generate image pairs that differ only in the embedded word while all other visual elements remain fixed. The model is trained to generate the benign image when conditioned on the original NSFW prompt. By training on a diverse set of NSFW and benign text-image pairs, the model learns to suppress NSFW text even for terms not seen during training.
@@
Finally, to evaluate the safety of vision generative models and equip the community with a reliable tool to monitor progress in this domain, we present ToxicBench, a comprehensive open-source benchmark built upon CreativeBench (Yang et al., 2024b). ToxicBench features a carefully curated dataset of textual prompts that trigger NSFW text generation, as well as the safety fine-tuning dataset used in our mitigation method. It also includes new metrics for text and image quality, and a robust evaluation pipeline. By exploring this novel threat vector and providing a standardized evaluation benchmark for the community, we aim to foster the development of safer multi-modal generative models. In summary, we make the following contributions:
@@
    We introduce a novel safety fine-tuning method that mitigates NSFW text in DMs by training on image pairs that differ only in the embedded text, where the NSFW term is replaced with a carefully chosen benign counterpart. The model is conditioned on the NSFW prompt but learns to generate the benign image, with LoRA updates applied only to localized text-generation layers. This setup enables the model to generalize suppression behavior to unseen NSFW terms while preserving both image and text quality.
@@ 2 Background and Related Work
Text-to-image Diffusion Models. DMs (Song and Ermon, 2020; Ho et al., 2020; Rombach et al., 2022) learn to approximate a data distribution by training a model, $\epsilon_{\theta}(x_{t},t,y)$, to denoise samples and reverse a stepwise diffusion process. Synthetic images are generated by initializing a sample with Gaussian noise, $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and iteratively subtracting the estimated noise at each time step $t=T,\ldots,1$, until a clean sample $x_{0}$ is reconstructed. Commonly, the denoising model $\epsilon_{\theta}(x_{t},t,y)$ is implemented using a U-Net (Ronneberger et al., 2015) (e.g., DeepFloyd IF) or transformer-based architectures (Vaswani, 2017) (e.g., SD3 (Esser et al., 2024)). Text-to-image DMs (Ramesh et al., 2022; Rombach et al., 2022; StabilityAI, 2023) include additional conditioning on some textual description $y$ in the form of a text embedding that is obtained by a pre-trained text encoder, such as CLIP (Radford et al., 2021) or T5 (Raffel et al., 2020). Initially, DMs failed to produce legible and coherent text within visuals, however, newer architectures, such as FLUX, Deep Floyd IF, SD3 and SDXL integrate multiple text encoders based on CLIP (Radford et al., 2021) or large language models like T5  (Raffel et al., 2020) that significantly improved the quality of the generated text.
@@
Layer-wise Control in Diffusion Models. Recent work has shown that specific layers in DMs are disproportionately responsible for rendering textual content within generated images (Staniszewski et al., 2025). These findings enable localized interventions that avoid full model fine-tuning, preserving general capabilities while modifying only the generative behavior tied to text rendering. We leverage this insight to fine-tune a small set of attention layers in each model family (e.g., joint attention in SD3 and cross-attention in SDXL and DeepFloyd IF) as part of our mitigation strategy.
@@
Harmful Visual Content Generation and Mitigation. Generative vision models have been shown to produce harmful content, such as NSFW imagery (Qu et al., 2023; Rando et al., 2022; Yang et al., 2024a), even when such content is not explicitly specified in prompts (Hao et al., 2024; Li et al., 2024). To detect this type of behavior, multiple dedicated detectors, e.g., (Berg, ; notAI tech, ) have been developed. Alternatively, large visual language model-based classifiers, relying, for example, on LLaVA (Liu et al., 2023), InstructBLIP (Dai et al., 2023), or GPT4V (OpenAI, ) have shown to be effective. Various mitigation techniques have been proposed. For instance, Safe Latent Diffusion (SLD) (Schramowski et al., 2023) guides generation away from unsafe concepts by adding a safety-conditioned loss during inference. Erase Stable Diffusion (ESD) (Gandikota et al., 2023) fine-tunes the model by steering the unconditional generation away from unsafe concepts using modified classifier-free guidance. Finally, Zong et al. (2024) build a safety-alignment dataset for fine-tuning vision language models. A complementary approach is explored by Safe-CLIP (Poppi et al., 2025), which targets the CLIP encoder underlying common DM architectures and performs multi-modal training that redirects inappropriate content while preserving embedding structure. However, these approaches are designed to address visual NSFW content (i.e., visual scenes of violence or nudity) and fail to tackle the issue of NSFW text embedded in the generated images as we show in Section 3.2, leaving this severe threat unaddressed.
@@
Harmful Text Generation and Mitigation. Large language models (LLMs) have been shown to generate NSFW text (Poppi et al., 2024; Gehman et al., 2020), despite safety alignment being in place (Wei et al., 2024; Ousidhoum et al., 2021). While NSFW text generation in LLMs involves discrete tokens, recent DMs rely on pretrained text encoders to condition image generation on natural language prompts. These encoders play a pivotal role in how textual information is translated into visual content. This shared reliance provides a technical basis for adapting safety interventions from the language domain to DMs. Most work in this domain focuses on fine-tuning the model to remove NSFW behavior, using either supervised examples (Adolphs et al., 2023) or reinforcement learning with human feedback (Ouyang et al., 2022; Bai et al., 2022). Other work operates on the neuron-level, identifies neurons that are responsible for toxic content and dampens these neurons (Suau et al., 2024). We evaluate the latest work on AURA (Suau et al., 2024) as a baseline and show that it suffers from the same limitations as existing solutions for the visual domain in preventing NSFW text embedding into images. This highlights the necessity of designing novel methods to address this threat in image generation.
@@ 3 Existing NSFW Solutions for Text or Vision Fail on Text in Images
The goal is to prevent the embedding of NSFW text in synthetic generated images. In this section, we explore naive solutions and existing baselines designed for the text or visual domains and show that they are ineffective in achieving this goal—either failing to prevent the generation of NSFW text or harming the model’s overall text generation ability significantly.
@@ 3.1 Naive Solutions Fail
As a very intuitive approach, one might want to treat the problem as purely text-based and attempt to solve it through the text prompt that causes the NSFW generation. This would involve an off-the-shelf toxicity detector, such as (Jigsaw, ; Hanu and Unitary team, 2020), to evaluate input prompts. NSFW prompts could then be rewritten with a language model before generation. However, this approach has multiple limitations. 1) First, whether certain words are perceived as NSFW depends on the visual context in the output. We observe that a variety of terms (e.g., Cocks or Penetrating) that can be perceived offensive without the right context, are not detected as NSFW by any off-the-shelf toxic text detectors we explored, e.g., (Hanu and Unitary team, 2020). For this reason, Hu et al. (2024) argue that effective NSFW filters need access to both input and output to avoid false negatives. In our case, although the input prompt may be classified as safe, the generated text in the output images can become offensive due to the contextual elements within the visual space. For instance, the word Penetrating, used in a cybersecurity setting, typically refers to the act of attacking a system. However, when presented in a different visual context, it may suggest a reference to a sexual act. 2) Classification-based toxicity detectors can overly restrict benign users and introduce latency. 3) Finally, this approach is restricted to API-based models with black box access but fails for open-source or locally deployed models, where users can simply bypass the re-writing step.
@@
Table 1: Harmful Content Detection. We assess the success of various NSFW detection approaches to identify images with embedded NSFW words. Multiheaded Detector (MHD(Qu et al., 2023) and the Stable Diffusion Filter (SD Filter(Rando et al., 2022) are solutions built for detecting NSFW visual scenes. OCR with Detoxify API (OCR+Detoxify) Hanu and Unitary team (2020) refers to our custom pipeline of using OCR to detect the words, and then performing NSFW classification with the Detoxify API. As a baseline, 100% of our NSFW words in the input prompt are classified as NSFW by Detoxify.
Attempt 2: Detecting and Censoring NSFW Text in Images. Alternatively, one could generate the image, locate the text, apply Optical Character Recognition (OCR) to extract it, classify the extracted text as NSFW or benign using a text-based toxicity detector, and then overwrite, blur, or censor NSFW text. While this approach shares all the limitations of the previous one (lack of context, latency, and non-applicability to open models), it has additional points of failure, namely the generation and the OCR. Already with small spelling errors or artifacts, the words are not correctly detected as NSFW anymore, even though still fully recognizable as offensive by a human observer. We quantify the detection success in the right column of Table 1 and plot examples of failure cases for NSFW detection in Figure 2. Overall, for FLUX—the model with the strongest text generation capabilities and, consequently, the highest OCR accuracy—this naive approach detects only 91% of NSFW samples, leaving 9% of harmful content undetected. Performance is even worse for other models, with detection rates dropping below 50% for SDXL. To explore whether visual NSFW detectors, i.e., the ones trained to detect NSFW visual scenes might be less easily fooled by the spelling mistakes, we also explore the detection success of two state-of-the-art vision detectors (Multiheaded Detector (Qu et al., 2023) and Stable Diffusion (SD) Filter (Rando et al., 2022)). The results in Table 1 show that these detectors fall even further behind the solution of combining OCR with text-based detection. SD Filter still achieves up to 46.45% detection accuracy for FLUX. This success rate is due to the underlying CLIP model, which enables the SD Filter to identify certain types of unsafe content even though it was not explicitly trained for text detection in images. CLIP’s ability to associate visual elements with textual descriptions contribute to this detection performance. Yet, with significant fractions of the NSFW samples undetected, and due to its conceptual limitations, this naive second attempt is also not sufficient to solve the problem.
Figure 2: OCR-based Detectors Insufficiency. We show SD3-generated images where the extracted text receives a low toxicity score (Hanu and Unitary team, 2020) ($<0.1$), while still being recognizable as offensive by human observers.

### v4

@@ Abstract
State-of-the-art Diffusion Models (DMs) produce highly realistic images. While prior work has successfully mitigated Not Safe For Work (NSFW) content in the visual domain, we identify a novel threat: the generation of NSFW text embedded within images. This includes offensive language, such as insults, racial slurs, and sexually explicit terms, posing significant risks to users. We show that all state-of-the-art DMs (e.g., SD3, SDXL, Flux, DeepFloyd IF) are vulnerable to this issue. Through extensive experiments, we demonstrate that existing mitigation techniques, effective for visual content, fail to prevent harmful text generation while substantially degrading benign text generation. As an initial step toward addressing this threat, we introduce a novel fine-tuning strategy that targets only the text-generation layers in DMs. Therefore, we construct a safety fine-tuning dataset by pairing each NSFW prompt with two images: one with the NSFW term, and another where that term is replaced with a carefully crafted benign alternative while leaving the image unchanged otherwise. By training on this dataset, the model learns to avoid generating harmful text while preserving benign content and overall image quality. Finally, to advance research in the area, we release ToxicBench, an open-source benchmark for evaluating NSFW text generation in images. It includes our curated fine-tuning dataset, a set of harmful prompts, new evaluation metrics, and a pipeline that assesses both NSFW-ness and text and image quality. Our benchmark aims to guide future efforts in mitigating NSFW text generation in text-to-image models, thereby contributing to their safe deployment.
@@ 1 Introduction
Warning: This paper contains examples of offensive language, including insults, and sexual or explicit terms, used solely for research and analysis purposes. State-of-the-art Diffusion Models (DMs) (Esser et al. 2024; StabilityAI 2023; Black Forest Labs 2024), have revolutionized the creation of realistic, detailed, and aesthetically impressive content. Despite their capabilities, these models often raise ethical and safety concerns, as they can inadvertently generate NSFW content, such as depictions of violence or nudity (Qu et al. 2023; Rando et al. 2022; Yang et al. 2024b).
@@
To mitigate the generation of NSFW content, prior work has focused extensively on addressing such issues in the visual space. Beyond the development of powerful NSFW detectors (Berg 2025; notAI tech 2025), these efforts, which include modifying training data (Zong et al. 2024), adding safety-based loss functions (Poppi et al. 2025; Gandikota et al. 2023), and steering generation to safe subspaces (Schramowski et al. 2023), have shown promising results in reducing explicit or harmful visual scenes.
@@
However, as DMs have grown more powerful, they now go beyond visual generation. In addition to generating realistic visuals, modern DMs now produce embedded text within images, such as captions, signs, or artistic typography (Esser et al. 2024; Chen et al. 2023; StabilityAI 2023; Black Forest Labs 2024). This advancement introduces a new challenge: as we show in Figure 1, all prominent state-of-the-art DMs, such as SD3 (Esser et al. 2024), Flux (Black Forest Labs 2024), DeepFloyd IF (StabilityAI 2023) and SDXL (Podell et al. 2023), can inadvertently produce NSFW or offensive text, such as explicit language or slurs that can be deeply offensive to viewers and raise significant ethical concerns. Even more, such text can escalate into more serious forms of toxic content, including targeted hate speech or ideologically charged propaganda, which makes their presence in generated images a nontrivial safety concern.
@@
We demonstrate that existing NSFW mitigation techniques (Gandikota et al. 2023; Poppi et al. 2025; Suau et al. 2024), while effective in addressing NSFW content in the visual or the language domain, are inadequate for handling embedded NSFW text in generated images without significantly degrading the models’ overall and (benign) text generations.
@@
Figure 1: Visual generative models output images with NSFW text. We evaluate 4 state-of-the-art DMs and observe that they easily generate NSFW text in the output images.
@@
As a first step toward addressing this threat, we introduce a novel method that performs lightweight fine-tuning on text-generation-relevant layers in DMs, previously identified by Staniszewski et al. (2025). By applying LoRA-based updates only to those layers, we enable efficient and focused mitigation. To supervise the intervention, we curate a safety fine-tuning dataset consisting of NSFW and benign prompt pairs that differ by a single word, where the harmful term is replaced with a carefully chosen benign counterpart. We generate image pairs that differ only in this embedded word while all other visual elements remain fixed. The model is trained to generate the benign image when conditioned on the original NSFW prompt. By training on a diverse set of NSFW and benign text-image pairs, the model learns to suppress NSFW text even for terms not seen during training.
@@
Importantly, unlike input or output filtering methods that are only effective in black-box scenarios where the model is accessed through an API, our approach modifies the model’s weights directly. This makes it applicable even in white-box or open-weight settings, where conventional filtering does not offer protection. Finally, to evaluate the safety of vision generative models and equip the community with a reliable tool to monitor progress in this domain, we present ToxicBench, a comprehensive open-source benchmark built upon CreativeBench (Yang et al. 2024a). ToxicBench features a carefully curated dataset of textual prompts that trigger NSFW text generation, as well as the safety fine-tuning dataset used in our mitigation method. It also includes new metrics for text and image quality, and a robust evaluation pipeline. By exploring this novel threat vector and providing a standardized evaluation benchmark for the community, we aim to foster the development of safer multi-modal generative models. In summary, we make the following contributions:
@@
    We introduce a novel safety fine-tuning method that mitigates NSFW text in DMs by training on image pairs that differ only in the embedded text, where the NSFW term is replaced with a carefully chosen benign counterpart. The model is conditioned on the NSFW prompt but learns to generate the benign image, with LoRA updates applied only to localized text-generation layers. This setup enables the model to generalize suppression behavior to unseen NSFW terms while preserving image and text quality.
@@ 2 Background and Related Work
DMs (Song and Ermon 2020; Ho, Jain, and Abbeel 2020; Rombach et al. 2022) learn to approximate a data distribution by training a model, $\epsilon_{\theta}(x_{t},t,y)$, to denoise samples and reverse a stepwise diffusion process. Synthetic images are generated by initializing a sample with Gaussian noise, $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and iteratively subtracting the estimated noise at each time step $t=T,\ldots,1$, until a clean sample $x_{0}$ is reconstructed. Commonly, the denoising model $\epsilon_{\theta}(x_{t},t,y)$ is implemented using a U-Net (Ronneberger, Fischer, and Brox 2015) (e.g., DeepFloyd IF) or transformer-based architectures (Vaswani 2017) (e.g., SD3 (Esser et al. 2024)). Text-to-image DMs (Ramesh et al. 2022; Rombach et al. 2022; StabilityAI 2023) include additional conditioning on some textual description $y$ in the form of a text embedding that is obtained by a pre-trained text encoder, such as CLIP (Radford et al. 2021) or T5 (Raffel et al. 2020). Initially, DMs failed to produce legible and coherent text within visuals, however, newer architectures, such as FLUX, Deep Floyd IF, SD3 and SDXL integrate multiple text encoders based on CLIP (Radford et al. 2021) or large language models like T5  (Raffel et al. 2020) that significantly improved the quality of the generated text.
Layer-wise Control in Diffusion Models.
@@
Recent work has shown that specific layers in DMs are disproportionately responsible for rendering textual content within generated images (Staniszewski et al. 2025). These findings enable localized interventions that avoid full model fine-tuning, preserving general capabilities while modifying only the generative behavior tied to text rendering. We leverage this insight to fine-tune a small set of attention layers in each model family (e.g., joint attention in SD3 and cross-attention in SDXL and DeepFloyd IF) as part of our mitigation strategy.
Harmful Visual Content Generation and Mitigation.
@@
Generative vision models have been shown to produce harmful content, such as NSFW imagery (Qu et al. 2023; Rando et al. 2022; Yang et al. 2024b), even when such content is not explicitly specified in prompts (Hao et al. 2024; Li et al. 2024). To detect this type of behavior, multiple dedicated detectors, e.g., (Berg 2025; notAI tech 2025) have been developed. Alternatively, large visual language model-based classifiers, relying, for example, on LLaVA (Liu et al. 2023), InstructBLIP (Dai et al. 2023), or GPT4V (OpenAI 2025) have shown to be effective. Various mitigation techniques have been proposed. For instance, Safe Latent Diffusion (SLD) (Schramowski et al. 2023) guides generation away from unsafe concepts by adding a safety-conditioned loss during inference. Erase Stable Diffusion (ESD) (Gandikota et al. 2023) fine-tunes the model by steering the unconditional generation away from unsafe concepts using modified classifier-free guidance. Finally, Zong et al. (2024) build a safety-alignment dataset for fine-tuning vision language models. A complementary approach is explored by Safe-CLIP (Poppi et al. 2025), which targets the CLIP encoder underlying common DM architectures and performs multi-modal training that redirects inappropriate content while preserving embedding structure. However, these approaches are designed to address visual NSFW content (i.e., visual scenes of violence or nudity) and fail to tackle the issue of NSFW text in the generated images as we show in Figure 2, leaving this severe threat unaddressed.
Harmful Text Generation and Mitigation.
@@
Large language models (LLMs) have been shown to generate NSFW text (Poppi et al. 2024; Gehman et al. 2020), despite safety alignment being in place (Wei et al. 2024; Ousidhoum et al. 2021). While NSFW text generation in LLMs involves discrete tokens, recent DMs rely on pretrained text encoders to condition image generation on natural language prompts. These encoders play a pivotal role in how textual information is translated into visual content. This shared reliance provides a technical basis for adapting safety interventions from the language domain to DMs. Most work in this domain focuses on fine-tuning the model to remove NSFW behavior, using either supervised examples (Adolphs et al. 2023) or reinforcement learning with human feedback (Ouyang et al. 2022; Bai et al. 2022). Other work operates on the neuron-level, identifies neurons that are responsible for toxic content and dampens these neurons. We evaluate the latest work (AURA) (Suau et al. 2024) as a baseline and show that it suffers from the same limitations as existing solutions for the visual domain in preventing NSFW text embedding into images. This highlights the necessity of designing novel methods to address this threat in image generation.
@@ 3 Existing NSFW Solutions for Text or Vision Fail on Text in Images
The goal is to prevent the embedding of NSFW text in synthetic generated images. In this section, we explore naive solutions and existing baselines designed for the text or visual domains and show their ineffectiveness in achieving this goal. They either fail to prevent the generation of NSFW text or harm the model’s text generation ability significantly.
@@ 3.1 Naive Solutions Fail
As a very intuitive approach, one might want to treat the problem as purely text-based and attempt to solve it through the text prompt that causes the NSFW generation. This would involve an off-the-shelf toxicity detector, such as (Jigsaw 2025; Hanu and Unitary team 2020), to evaluate input prompts. NSFW prompts could then be rewritten with a language model before generation. However, this approach has multiple limitations. 1) First, whether certain words are perceived as NSFW depends on the visual context in the output. We observe that a variety of terms (e.g., Cocks or Penetrating) that can be perceived offensive without the right context, are not detected as NSFW by any off-the-shelf toxic text detectors we explored, e.g., (Hanu and Unitary team 2020). For this reason, Hu et al. (2024) argue that effective NSFW filters need access to both input and output to avoid false negatives. In our case, although the input prompt may be classified as safe, the generated text in the output images can become offensive due to the contextual elements within the visual space.
For instance, the word Penetrating, used in a cybersecurity setting, typically refers to the act of attacking a system. However, when presented in a different visual context, it may suggest a reference to a sexual act. 2) Classification-based toxicity detectors can overly restrict benign users and introduce latency. 3) Finally, and most importantly, this approach is restricted to API-based models with black-box access but fails for open-source or locally deployed models, where users can simply bypass the re-writing step. In contrast, our solution directly modifies the model’s weights and is therefore applicable in white-box settings, including open-weight models that users can run locally.
Attempt 2: Detecting and Censoring NSFW Text in Images.
Alternatively, one could generate the image, locate the text, apply Optical Character Recognition (OCR) to extract it, classify the extracted text as NSFW or benign using a text-based toxicity detector, and then overwrite, blur, or censor NSFW text. While this approach shares all the limitations of the previous one (lack of context, latency, and non-applicability to open models), it has additional points of failure, namely the generation and the OCR. Already with small spelling errors or artifacts, the words are not correctly detected as NSFW anymore, even though still fully recognizable as offensive by a human observer.
Figure 2: OCR-based Detectors Insufficiency. We show SD3-generated images where the extracted text receives a low toxicity score (Hanu and Unitary team 2020) (<0.1<0.1), while still being recognizable as offensive by human observers.
@@
Table 1: Harmful Content Detection. We assess the success of various NSFW detection approaches to identify images with embedded NSFW words. Multiheaded Detector (MHD(Qu et al. 2023) and the Stable Diffusion Filter (SD Filter(Rando et al. 2022) are solutions built for detecting NSFW visual scenes. OCR with Detoxify API (OCR+Detoxify) (Hanu and Unitary team 2020) refers to our custom pipeline of using OCR to detect the words, and then performing NSFW classification with the Detoxify API. As a baseline, 100% of our NSFW words in the input prompt are classified as NSFW by Detoxify.
We quantify the detection success in the right column of Table 1 and plot examples of failure cases for NSFW detection in Figure 2. Overall, for FLUX this naive approach detects only 91% of NSFW samples, leaving 9% of harmful content undetected. Performance is even worse for other models, with detection rates dropping below 50% for SDXL. To explore whether visual NSFW detectors, i.e., the ones trained to detect NSFW visual scenes might be less easily fooled by the spelling mistakes, we also explore the detection success of two state-of-the-art vision detectors (Multiheaded Detector (Qu et al. 2023) and Stable Diffusion (SD) Filter (Rando et al. 2022)). The results in Table 1 show that these detectors fall even further behind the solution of combining OCR with text-based detection. SD Filter still achieves up to 46.45% detection accuracy for FLUX. This success rate is due to the underlying CLIP model, which enables the SD Filter to identify certain types of unsafe content even though it was not explicitly trained for text detection in images. CLIP’s ability to associate visual elements with textual descriptions contribute to this detection performance. Yet, with significant fractions of the NSFW samples undetected, and due to its conceptual limitations, this naive second attempt is also not sufficient to solve the problem.